xx NullPsi4IO

NullPsi4IO

Maria Babiuc <babiuc@marshall.edu>

October 21 2009

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11maskxtype 10Scope: privatex01 KEYWORD

11Description: type of mask for error in J

11 Range 10 x01 Default: none
11none  10no mask
11EGxmask  10mask all the active points
11EVxmask  10mask only the evolution points
11EQxmask  10mask only the equator points

11

11nullpsiintxmaxevolvedarraysize 10Scope: restrictedx01  INT

11Description: The maximum size of evolved grid arrays used by NullPsiInt

11 Range 10 x01 Default: 1
111:*  10The size of the generator grid arrays

11

11nullpsiintxnumarrayevolvedvars 10Scope: restrictedx01  INT

11Description: The maximum number of evolved grid arrays used by NullPsiInt

11 Range 10 x01 Default: 4
114:4  10Is is the real and imaginary part of the RHS (Psi) and the News obtained by integrating Psi

11

11nxangxevxoutsidexeq  10Scope: shared from NULLGRIDx01INT

11

11nxangxghostxpts  10Scope: shared from NULLGRIDx01INT

11

11nxangxptsxinsidexeq  10Scope: shared from NULLGRIDx01INT

11

11nxangxstencilxsize  10Scope: shared from NULLGRIDx01INT

11

11nxradialxpts  10Scope: shared from NULLGRIDx01INT

11

7 Interfaces General Implements: nullpsiint

Inherits: nullvars nullgrid nullnews nullinterp nullevolve

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 rexdotnewsb 10 rexdotNewsB x01  compact110
11                 10  x01 description11real dotNews for MoL - this is the real part of Psi4 (RHS)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 imxdotnewsb 10 imxdotNewsB x01  compact110
11                 10  x01 description11imag dotNews for MoL - this is the imaginary part of Psi4 (RHS)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 rexpsiint 10 rexPsiInt x01 compact110
11                 10  x01 description11real integrated Psi for MoL - this is the real part of NewsBxPsi (RHS)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 imxpsiint 10 imxPsiInt x01 compact110
11                 10  x01 description11imag integrated Psi for MoL - this is the imaginary part of NewsBxPsi4 (RHS)
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11

7.0.2 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 newsbxmask 10 NewsBxmask x01  compact110
11                 10  x01 description11News inside equator
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 psi4xmask 10 Psi4xmask x01 compact110
11                 10  x01 description11Psi4 inside equator
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 newsbxpsi 10 NewsBxPsi x01 compact110
11                 10  x01 description11integrated Psi
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullPsiInt. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11NewsBxmask Psi4xmask NewsBxPsi rexdotNewsB[2] imxdotNewsB[2] rexPsiInt[2] imxPsiInt[2]11x

11x  11x

Scheduled Functions

NullEvolxInitial x nullpsiintxinitmol x initial data x

11x11After: 11nullevolxinitialslice
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

MoLxRegister x nullpsiintxmolregister x register variables for mol x

11x11Language:11c
11 x 11Type: 11function

11

MoLxCalcRHS x nullpsiintxdotnewspsi x calculate the integral of psi4 x

11x11Language:11fortran
11 x 11Sync:  11newsbxpsi
11 x 11Type: 11function